# Scalable Oversight via Partitioned Human Supervision

### v1

@@ Appendix F Estimator Configurations
    Complementary-label estimator with sample size fixed at (K1)(K-1) times the ordinary case, i.e. $n_{\mathrm{c}}=(K-1)n_{\mathrm{o}}$. Following the protocol in Appendix B, each instance with oracle label $y$ is converted into (K1)(K-1) complementary “No” labels $\{\bar{y}\in\mathcal{Y}\setminus\{y\}\}$, and the ordinary label is discarded; thus no ordinary labels are present in this setting.
@@
    Maximum-likelihood estimator obtained by solving the joint likelihood equation using both ordinary and complementary samples. Asymptotically equivalent to IVW but involving a nonlinear update (Appendix C).
@@
Appendix G Implementation details on ADAS and AFlow
To ensure a fair comparison, we constructed validation and test splits for each benchmark, following the numbers and ratios used in Hu et al. (2025). Specifically, we sampled 120 validation and 800 test examples for Math-MC and Medical Abstract, and 88 validation and 458 test examples for GPQA. For AFlow, we set the number of validation runs to 3 and used the same validation set as ADAS, rather than the separate validation subset reported in Zhang et al. (2025c).
For code and prompts, we largely follow the official implementations of Hu et al. (2025) and Zhang et al. (2025c), with two minor modifications for reproducibility in our setting.
@@ Appendix G Implementation details on ADAS and AFlow
Appendix H Estimator Variance Analysis

### v2

@@ Appendix I Estimator Configurations
    Complementary-label estimator with sample size fixed at (K1)(K-1) times the ordinary case, i.e. $n_{\mathrm{c}}=(K-1)n_{\mathrm{o}}$. Following the protocol in Appendix C, each instance with oracle label $y$ is converted into (K1)(K-1) complementary “No” labels $\{\bar{y}\in\mathcal{Y}\setminus\{y\}\}$, and the ordinary label is discarded; thus no ordinary labels are present in this setting.
@@
    Maximum-likelihood estimator obtained by solving the joint likelihood equation using both ordinary and complementary samples. Asymptotically equivalent to IVW but involving a nonlinear update (Appendix D).
@@
Appendix J Implementation details on ADAS and AFlow
To ensure a fair comparison, we constructed validation and test splits for each benchmark, following the numbers and ratios used in Hu et al. (2025). Specifically, we sampled 120 validation and 800 test examples for Math-MC and Medical Abstract, and 88 validation and 458 test examples for GPQA. For AFlow, we set the number of validation runs to 3 and used the same validation set as ADAS, rather than the separate validation subset reported in Zhang et al. (2025b).
For code and prompts, we largely follow the official implementations of Hu et al. (2025) and Zhang et al. (2025b), with two minor modifications for reproducibility in our setting.
@@
Appendix K Estimator Variance Analysis
@@
Appendix L Deviation from the Ord-Eval Reference
For completeness, Table 3 reports the absolute deviation ($\Delta$) between each estimator and the Ord-Eval oracle on each benchmark. Column-wise minima are shown in bold. These values show that mixture estimators (IVW and ML) reduce the gap to the Ord-Eval reference compared with both ordinary-only and complementary-only estimators across most benchmarks, indicating that they provide more robust estimates. Ord and Comp-$n_{\mathrm{o}}$ can occasionally produce point estimates that are numerically closer to Ord-Eval (as in MedQA and GPQA). However, as discussed in Section 3.1, their standard errors are substantially larger, suggesting poorer reliability and limited reproducibility. This further supports the robustness of the mixture estimators.
Table 3: Absolute deviation ($\Delta$; in percentage points) from the Ord-Eval oracle. Column-wise minima are shown in bold.
Estimator MMLU-Pro MedQA-USMLE GPQA MATH MATH(CoT) Avg $\Delta$
Ord 0.36 0.23 4.65 3.35 1.00 1.92
Comp-$n_{o}$ 0.97 0.01 0.35 4.23 3.45 1.80
Comp-Var 2.30 2.05 4.15 3.11 2.54 2.83
IVW-0.5 0.08 1.05 5.76 0.77 0.33 1.60
IVW 0.00 0.80 5.62 0.66 0.03 1.42
ML 0.03 1.01 5.59 0.54 0.24 1.48
Appendix M Case Study: Behavior of Agents Trained with Partial Feedback
In this section, we visualize and analyze the internal mechanisms of the agentic systems produced by ADAS and AFlow on the Medical Abstracts dataset. This case study illustrates how agentic workflows learned under partial feedback acquire coherent, structured decision processes. Our goal is to show that, even with weak supervision, these systems diverge meaningfully and improve from fixed workflows.
(a) Principled Medical Abstract Classifier (ADAS)
(b) Review-Generate–Ensemble Workflow (Aflow)
Figure 5: Workflows of the best ADAS and AFlow agents learned under partial-feedback training. Complementary label rewards lead ADAS to adopt a principle-extraction–guided classifier, while AFlow learns a review-generate–ensemble pipeline.
Figure 5 presents the best-performing agentic workflows discovered by ADAS and AFlow. The ADAS system, which we refer to as the Principled Medical Abstract Classifier, and the AFlow workflow, which name was chosen for clarity, both adopt multi-stage decision pipelines rather than directly producing a single-shot answer, as done by baseline methods.
In Figure 5(a), ADAS performs a consistent two-stage decomposition. Instead of mapping an abstract directly to a disease label, the agent first produces a numbered list of key medical principles or mechanistic factors extracted from the text. These principles are then passed to a separate classifier, which uses them as evidence for the final prediction. Figure 5(b) shows that AFlow adopts a multi-candidate reasoning pipeline. The agent first generates a structured review of the abstract using a task-specific prompt. Conditioned on this shared reasoning trace, it then produces several candidate answers. A final ensemble operator aggregates these candidates and selects the most consistent prediction.
Such modular, review-based, and multi-candidate workflows do not appear in established baselines, which rely on fixed or single-pass reasoning. The emergence of these structures demonstrates that complementary label supervision can shape the agent’s decision process, not merely its accuracy. Partial feedback encourages the agent to construct interpretable intermediate representations and use them as evidence, even without full supervision. This confirms that complementary label training provides a behavior shaping signal that yields more structured and interpretable workflows.
Appendix N Large Language Models Usage Statement
LLMs assisted us in formulating our ideas, working on extensions and analysis, and polishing the writing. All of the LLM outputs were subsequently verified, revised and modified by the authors. Any errors or inappropriate statements remain the responsibility of the authors.